OpenStudy (jessicawade):

Graphs of rational functions help! will medal 2 questions

OpenStudy (jessicawade):

need help with these questions! can you help me? 1. Write a specific formula to describe the variation: x varies inversely with the square of y; x = 6 when y = 4. 8.Write a general formula to describe the variation: x varies inversely with the cube of y.

OpenStudy (dumbcow):

inverse variation means as one thing goes up the other goes down the functions will be fractions, one will be in the denominator 1) let y^2 be in denominator \[x = \frac{k}{y^2}\] solve for "k" using given point 6 = k/16 ---> k = 96

OpenStudy (jessicawade):

ok

OpenStudy (jessicawade):

\[x= -96/ \sqrt y\]

OpenStudy (jessicawade):

wait wrong

OpenStudy (jessicawade):

\[x= 96 \over y^2\]

OpenStudy (jessicawade):

@dumbcow

OpenStudy (dumbcow):

yes correct

OpenStudy (jessicawade):

Thanks :)

OpenStudy (dumbcow):

for 8) the general formula just leave the unknown constant "k" x = k/y^3

OpenStudy (jessicawade):

ohhh so its x= k/y^3?

OpenStudy (dumbcow):

for the 2nd question yes

OpenStudy (jessicawade):

yay! thanks

OpenStudy (jessicawade):

i get the words mixed up sometimes but 8 was obvious -_-

OpenStudy (dumbcow):

haha :)

OpenStudy (jessicawade):

ok i have 1 more question if that is ok

OpenStudy (jessicawade):

Suppose that the time it takes to do a job is inversely proportional to the number of workers. That is, the more workers on the job the less time required to complete the job. If it takes 3 workers 16 days to finish a job, how long will it take 8 workers?

OpenStudy (jessicawade):

@dumbcow